• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M Defurne

    Log entry time 07:56:51 on May 05, 2011

    Entry number 358248

    keyword=Owl shift summary

    >>Shift Crew

    SL: Maxime Defurne TO: Paula Aguilera 3rd: Franco Garibaldi

    >>Shift Summary

    00:00 Shift starts with run 4177 for kinematics 4.1 with dummy 20cm. 00:50 Move target from dummy 20cm to dummy 10cm for production. 02:20 Change kinematics 4.1 to 5.05 03:10 Start first production run 4181 on He3 20cm for kinematics 5.05 with only the L HRS. Indeed we still need to set up the R Q3. 03:30 Run 4181 stops because current in right dipole is correct but the NMR is not locked. We will start a new run to take data on both spectrometers. 03:40 Coda crashed for run 4182 and 4183 04:13 The Right NMR is not locked yet. Moreover a red alarm appears because it doesn't work. As the current is stable, we will take data. 04:15 Coda just crashed. Runs 4185 and 4186 are junked. 05:15 We can't make a replay of the first 50000 events of the run 4188. 06:00 We changed the target from He3 20cm to Dummy 20cm. 06:12 run 4190 , prescales are not correct!!!!! 07:00 No more beam for beam studies

    >>Run Summary --KINEMATICS 4.1-- run Type Target Current Good Events Accumulated 4177 MISC Dummy 20cm 40 uA 43 min 4178 Prod Dummy 10cm 40 uA 312k 312k 4179 Prod Dummy 10cm 40 uA 344k 656k 4180 Prod Dummy 10cm 40 uA 190k 846k

    --Kinematics 5.05--

    run Type Target Current Good Events 4181 Prod He3 20cm 120 uA 4182 JUNK 4183 JUNK 4184 Prod He3 20cm 120 uA 4185 JUNK 4186 JUNK 4187 Prod He3 20cm 120 uA 4188 Prod He3 20cm 120 uA 4189 Prod He3 20cm 120 uA 4190 PRESCALES ARE NOT GOOD!!!! 4191 MISC dummy 20cm 40 uA 25 min 4192 MISC dummy 20cm 40 uA 25 min

    Sorry if there isn't the number of good events, we didn't really understand how to calculate it. Good runs for He3 kin5.05 are 4181(left arm only),4184,4187,4188